OpenStudy (anonymous):

The length of a rectangle exceeds the width by 17 and the area=18 what is the length and the width?

rishavraj (rishavraj):

let the width be x then length is (x + 17) nd area = l*b

OpenStudy (anonymous):

so 4x+34=4.25 Is that correct?

rishavraj (rishavraj):

no its suppposed to be x*(x + 17) =18...solve for x

OpenStudy (anonymous):

So x*x + 17x= 18 right solve for x and x=-1 and x=18 how is that correct??? Plug those into the original equation and it doesnt work???

OpenStudy (anonymous):

I have done this equation several times and the solutions are not getting me the area=18? so what am I doing wrong????

rishavraj (rishavraj):

see x^2 + 17x - 18 =0 x^2 - x +18x - 18 = 0 x(x - 1) + 18(x - 1) = 0 (x -1)(x + 18) =0 so x = 1 & -18...ignore -18

OpenStudy (anonymous):

Oh geeze I had the wrong sign. I knew I was doing something wrong. Thank you! that was driving me crazy!!! Have a wonderful day!

rishavraj (rishavraj):

i will surely have....